OpenStudy (anonymous):

Write y = 2/3 x + 3 in standard form using integers.

OpenStudy (hoblos):

first multiply by 3 3y = 2x +9 2x-3y +9 =0

OpenStudy (anonymous):

is that answer because theres 4 choices A.–2x + 3y = 9 B.–2x + 3y = 3 C.3x – 2y = 9 D.–2x – 3y = 9

OpenStudy (hoblos):

2x-3y +9 =0 2x-3y = -9 multiply by -1 –2x + 3y = 9 so A

OpenStudy (anonymous):

thank you

OpenStudy (hoblos):

my pleasure XD